OpenStudy (lexy148):

how do i solve, If 160lb of an alloy contains 76.8lb of copper, what peercent of the alloy is copper?

OpenStudy (lexy148):

can you explain how you got that answer

hero (hero):

@cj49, giving direct answers on OS is forbidden. Guiding, explaining, demonstrating, and tutoring is welcome here.

OpenStudy (cj49):

76.81/160x100=48

OpenStudy (cj49):

ill be more careful from next time@Hero

hero (hero):

You can use this equation for problems like these: \[\frac{\text{part}}{\text{whole}} = \frac{\%}{100}\] Then input the given values. In this case \(\%\) is \(x\) since it is unknown: \[\frac{76.8}{160} = \frac{x}{100}\]

OpenStudy (lexy148):

oh okay thank you i get it now @Hero

hero (hero):

You're welcome